OpenStudy (anonymous):

can someone please explain EACH step that are necessary to Solve 4x-2(x+3)=x+1

OpenStudy (saifoo.khan):

4x-2(x+3)=x+1

OpenStudy (saifoo.khan):

4x -2x -6 =x+1

OpenStudy (saifoo.khan):

2x -6 = x+1 2x - x = 6+1 x = 7

OpenStudy (anonymous):

distribution, simplification, using addition property of equality, simplification

OpenStudy (anonymous):

can you explain how you got the answer please